Getting lost in completing the square for 2x^2+5x+11=0
I am able to get to 2x^2+5x=-11 then do I divide everything by 2? That would give me a really complicated fraction of x^2+5x/2=-11/2 then I am lost...
Melody

Answer by stanbon(75887)

completing the square for
2x^2+5x+11=0
---
2x^2 + 5x = -11
2(x^2 + (5/2)x + (5/4)^2 = -11 + 2*(5/4)^2
----
2(x+(5/4))^2 = (-176/16)+(50/16)
2(x+(5/4))^2 = (-126/16)
(x+(5/4))^2 = (-63/16)
x+(5/4) = (3/4)sqrt(7)i or x+(5/4) = -(3/4)sqrt(7)i
---
x = -(5/4)+(3/4)sqrt(7)i or x = -(5/4)-(3/4)sqrt(7)i
=========================
Cheers,
Stan H.